# Calibrating Deep Neural Network using Euclidean Distance

### v1

@@ 3.1. Focal Calibration Loss
Our observations in Figure 3 indicate that focal loss tends to produce higher-entropy predictions, resulting in outputs closer to uniform probabilities (e.g., 0.5 in binary classification). This occurs even when the true probability $\eta_{k}(\bm{x})$ is near 0 or 1 (i.e., for easy samples), leading to underconfident predictions and reduced calibration reliability. Charoenphakdee et al. (Charoenphakdee et al., 2021) also demonstrated that focal loss is classification-calibrated but not strictly proper. To address calibration issues, we propose the Focal Calibration Loss (FCL) by adding a proper loss term to the focal loss.
@@
Figure 3. Binary classification ($K=2$): Optimal predicted probabilities $\hat{p}$ for ground-truth probabilities $q\in[0,1]$ using Cross-Entropy Loss, Focal Loss, and Focal Calibration Loss.
[…]
tbc
@@
Let $f:\mathcal{X}\rightarrow\Delta^{K}$ be a classifier mapping inputs $\bm{x}_{i}\in\mathcal{X}$ to predicted probabilities $\hat{\bm{p}}_{i}=f(\bm{x}_{i})=[\hat{p}_{i1},\hat{p}_{i2},\ldots,\hat{p}_{iK}]^{\top}\in\Delta^{K}$. Let $\bm{y}_{i}=[y_{i1},y_{i2},\ldots,y_{iK}]^{\top}\in\Delta^{K}$ be the one-hot encoded true labels. We define the Focal Calibration Loss $\mathcal{L}_{\text{FCL}}^{\gamma,\lambda}$ as:
@@
(11) $\mathcal{L}_{\text{FCL}}^{\gamma,\lambda}=\frac{1}{N}\sum_{i=1}^{N}\left[\mathcal{L}_{\text{focal}}(f(\bm{x}_{i}),\bm{y}_{i})+\lambda\,\mathcal{L}_{\text{calib}}(f(\bm{x}_{i}),\bm{y}_{i})\right],$
@@
where $\gamma\geq 0$ and $\lambda\geq 0$ are hyperparameters controlling the focus and calibration trade-off, respectively.
(12) $\mathcal{L}_{\text{focal}}(f(\bm{x}_{i}),\bm{y}_{i})$ $=-\sum_{k=1}^{K}y_{ik}\,(1-\hat{p}_{ik}(\bm{x}_{i}))^{\gamma}\,\log\hat{p}_{ik}(\bm{x}_{i}),$
@@
(13) $\mathcal{L}_{\text{calib}}(f(\bm{x}_{i}),\bm{y}_{i})$ $=\left\|f(\bm{x}_{i})-\bm{y}_{i}\right\|_{2}^{2}=\sum_{k=1}^{K}\left(\hat{p}_{ik}(\bm{x}_{i})-y_{ik}\right)^{2}.$
By directly optimizing calibration, we minimize the calibration error by reducing the Euclidean norm between $\hat{\bm{p}}_{i}$ and $\bm{y}_{i}$, even without access to the true class-posterior probabilities ($\bm{\eta}$). Consequently, FCL simultaneously constrains the bounds on overconfidence and underconfidence (Thm. 1) while ensuring classification calibrated with the strict propriety (Thm. 7), and achieves a smaller or equal post-processing gap compared to traditional focal loss (Thm. 8), thereby leading to better-calibrated models.
4. Theoretical Evidence
4.1. Mitigate OC/UC Beyond Focal Loss
Our goal is to show that minimizing the FCL leads to predicted probabilities $\hat{\bm{p}}(\bm{x})$ that are closer to $\bm{\eta}(\bm{x})$, thus mitigating OC/UC.
Theorem 1.
Let $f$ be a classifier trained by minimizing the Focal Calibration Loss $\mathcal{L}_{\text{FCL}}^{\gamma,\lambda}$ with $\lambda>0$. Then, for any $\bm{x}\in\mathcal{X}$, the overconfidence and underconfidence are bounded by the Euclidean distance between the predicted probabilities and the true class-posterior probabilities:
(14) $\left|\max_{k}\hat{p}_{k}(\bm{x})-\max_{k}\eta_{k}(\bm{x})\right|\leq\left\|\hat{\bm{p}}(\bm{x})-\bm{\eta}(\bm{x})\right\|_{2}.$
Therefore, minimizing $\mathcal{L}_{\text{FCL}}^{\gamma,\lambda}$ reduces overconfidence and underconfidence by directly minimizing $\left\|\hat{\bm{p}}(\bm{x})-\bm{\eta}(\bm{x})\right\|_{2}^{2}$. Since the Euclidean norm aggregates the squared differences across all classes, it provides an upper bound on any individual difference (Appendix A.2 and A.3).
4.2. Classification-Calibrated
In this section, we theoretically prove that minimizing the focal calibration risk $\mathcal{R}^{\mathcal{L}_{\text{FCL}}^{\gamma,\lambda}}$ yields the Bayes-optimal classifier, which ensures the highest possible expected accuracy in classification tasks (Zhang, 2004). To demonstrate this, we show that the Focal Calibration Loss (FCL) is classification-calibrated (Bartlett et al., 2006; Tewari and Bartlett, 2007).
Definition 2.
(Pointwise Conditional Risk). The pointwise conditional risk $\mathcal{R}^{\mathcal{L}}$ of an input $\bm{x}$ with its class-posterior probability $\bm{\eta}(\bm{x})$ is defined as:
(15) $\mathcal{R}^{\mathcal{L}}(\bm{q}(\bm{x});\bm{\eta}(\bm{x}))=\sum_{y\in\mathcal{Y}}\eta_{y}(\bm{x})\mathcal{L}(\bm{q}(\bm{x}),\bm{e}_{y}).$
Intuitively, the pointwise conditional risk $\mathcal{R}^{\mathcal{L}}$ corresponds to the expected penalty for a data point $\bm{x}$ when using $\bm{q}(\bm{x})$ as a score function. $\mathcal{L}(\bm{q}(\bm{x}),\bm{e}_{y})$ denotes the loss incurred by predicting $\bm{q}(\bm{x})$ when the true class is $y$, where $\bm{e}_{y}\in\{0,1\}^{K}$ is a one-hot encoded vector corresponding to class $y$.
Definition 3.
(Classification-Calibrated Loss (Bartlett et al., 2006; Tewari and Bartlett, 2007)). Consider a surrogate loss $\mathcal{L}$. Let $\bm{q}^{\mathcal{L},*}=\arg\min_{\bm{q}}\mathcal{R}^{\mathcal{L}}(\bm{q}(\bm{x});\bm{\eta}(\bm{x}))$ be the minimizer of the pointwise conditional risk. $\mathcal{L}$ is classification-calibrated, if $\mathcal{R}^{\mathcal{L}_{0\text{-}1}}(f^{\bm{q}^{\mathcal{L},*}})=\mathcal{R}^{\mathcal{L}_{0\text{-}1}}(f^{\mathcal{L}_{0\text{-}1},*})$.
Classification calibration guarantees that the minimizer of the pointwise conditional risk of a surrogate loss will yield the Bayes-optimal classifier. This definition suggests that by minimizing a classification-calibrated loss, even if $\bm{q}^{\mathcal{L},*}(\bm{x})$ is not equal to the true class-posterior probability $\eta(\bm{x})$, we can still achieve the Bayes-optimal classifier from $\bm{q}^{\mathcal{L},*}(\bm{x})$ as long as their decision rules match.
Definition 4.
(Strictly Order-Preserving Property). Let $\mathcal{X}$ be the input space, and $\mathcal{Y}=\{1,2,\ldots,K\}$ be the label space with $K$ classes. For an input sample $\bm{x}\in\mathcal{X}$, let $\bm{q}=[q_{1}(\bm{x}),q_{2}(\bm{x}),\ldots,q_{K}(\bm{x})]^{\top}$ denote the predicted probability distribution over the $K$ classes, and $\bm{\eta}(\bm{x})=[\eta_{1}(\bm{x}),\eta_{2}(\bm{x}),\ldots,\eta_{K}(\bm{x})]^{\top}$ denote the true class-posterior probability distribution. Let $\mathcal{L}(\bm{q},y)$ be a loss function, where $\bm{q}$ is the predicted probability distribution and $y$ is the true label. Formally, let $q_{i}^{\mathcal{L},*}(\bm{x})$ denote the minimizer of the pointwise conditional risk $\mathcal{R}^{\mathcal{L}}(\bm{x})$. The loss function $\mathcal{L}$ is classification-calibrated if and only if:
(16) $q_{i}^{\mathcal{L},*}(\bm{x})<q_{j}^{\mathcal{L},*}(\bm{x})\Rightarrow\eta_{i}(\bm{x})<\eta_{j}(\bm{x}).$
for all $i,j\in\{1,2,\ldots,K\}$ and for all $\bm{x}\in\mathcal{X}$. This ensures that the predicted probabilities’ order matches the true probabilities’ order, leading to a Bayes-optimal classifier.
A loss function is said to be classification-calibrated if it satisfies the strictly order-preserving property. This property ensures that the minimizer of the pointwise conditional risk preserves the order of the true class-posterior probabilities. To simplify notation, we use $\bm{q}^{\gamma,\lambda,*}$ to denote $\bm{q}^{{\mathcal{L}_{\text{FCL}}^{\gamma,\lambda}},*}$, i.e., the focal calibration risk minimizer with parameters $\gamma$ and $\lambda$. Since $\bm{q}^{\gamma,\lambda,*}$ preserves the order of $\eta$, it follows that $\arg\max_{y}\bm{q}^{\gamma,\lambda,*}_{y}(\bm{x})=\arg\max_{y}\eta_{y}(\bm{x})$. Thus, the Bayes-optimal classifier can be achieved by minimizing the focal calibration risk minimizer, i.e., $\mathcal{R}^{\mathcal{L}_{0\text{-}1}}(f^{\bm{q}^{\gamma,\lambda,*}})=\mathcal{R}^{\mathcal{L}_{0\text{-}1}}(f^{\mathcal{L}_{{0\text{-}1},*}})$. Our proof demonstrates that the FCL possesses the strictly order-preserving property (Appendix A.5), which is sufficient for classification calibration (Zhang, 2004). The following theorem guarantees that the focal calibration loss is classification-calibrated.
4.3. Strictly Proper
In this section, we examine the FCL in the context of class-posterior probability estimation. We theoretically demonstrate that the simplex output of the focal calibration risk minimizer $\bm{q}^{\gamma,\lambda,*}$ does not produce the true class-posterior probability. To be suitable for class-posterior probability estimation, a surrogate loss must be strictly proper, defined as follows.
Definition 5.
(Strictly Proper Loss (Buja et al., 2005; Gneiting and Raftery, 2007; Shuford Jr et al., 1966)). We say that a loss $\mathcal{L}:\Delta^{K}\times\Delta^{K}\rightarrow\mathbb{R}$ is strictly proper if $\mathcal{L}(\hat{\bm{p}},\bm{y})$ is minimized iff $\hat{\bm{p}}=\bm{y}$.
Strict properness can be viewed as a fundamental requirement for a loss function when estimating the true class-posterior probability (Williamson et al., 2016). When comparing the ground truth probability $\bm{y}$ and its estimate $\hat{\bm{p}}$, we want the loss function to be minimized if and only if $\hat{\bm{p}}=\bm{y}$, indicating that the probability estimation is correct. Strict properness is a more stringent requirement than classification calibration since all strictly proper losses are classification-calibrated; however, the reverse does not hold true (Reid and Williamson, 2010; Williamson et al., 2016). While focal loss has been proven to be not strictly proper, we demonstrate that calibration loss is strictly proper in Thm. 6 (Appendix A.7), and focal calibration loss is strictly proper in Thm. 7 (Appendix A.8).
Theorem 6.
The calibration loss defined as the mean squared $L_{2}$ norm, $\mathcal{L}_{\text{calib}}(f(\bm{x}_{i}),\bm{y}_{i})=\left\|f(\bm{x}_{i})-\bm{y}_{i}\right\|_{2}^{2}=\sum_{k=1}^{K}\left(\hat{p}_{ik}-y_{ik}\right)^{2},$ is a strictly proper scoring rule for both binary and multiclass classification.
Theorem 7.
The Focal Calibration Loss $\mathcal{L}_{\text{FCL}}^{\gamma,\lambda}$ is strictly proper iff $\lambda>0$ and $\gamma\geq 0$.
4.4. Reduction of Post-Processing Gap
Theorem 8.
@@ 4.4. Reduction of Post-Processing Gap
For any predictor $f:\mathcal{X}\to[0,1]$ and any distribution $\mathcal{D}$ over $\mathcal{X}\times\{0,1\}$, the focal calibration loss $\mathcal{L}_{\text{FCL}}^{\gamma,\lambda}$ results in a smaller or equal post-processing gap compared to the focal loss:
(17) $\text{pGap}_{\mathcal{D}}(\mathcal{L}_{\text{FCL}})\leq\text{pGap}_{\mathcal{D}}(\mathcal{L}_{\text{focal}}).$
Thm. 8 indicates that incorporating the calibration loss term into the FCL objective function effectively reduces the potential improvement achievable through post-processing. This suggests that models trained with FCL are inherently better calibrated, as it reduces the calibration error (Appendix A.14).
5. Experimental Results

### v2

@@ 3.1 Focal Calibration Loss
Our observations in Fig. 2 suggest that focal loss often yields higher-entropy predictions, leading to underconfident outputs even when the true probability $\eta_{k}(\boldsymbol{x})$ is near 0 or 1 (i.e., for “easy” samples). Such underconfidence undermines calibration reliability. Although Charoenphakdee et al. (Charoenphakdee et al., 2021) proved that focal loss is classification-calibrated, it is not strictly proper. We address these calibration issues by proposing the Focal Calibration Loss (FCL), which appends a proper loss term to focal loss.
@@
Figure 2: Binary classification ($K=2$): Optimal predicted probabilities $\hat{p}$ for ground-truth probabilities $q\in[0,1]$ using Cross-Entropy Loss, Focal Loss, and Focal Calibration Loss.
Definition 3.1 (Focal Calibration Loss (FCL)).
@@
Let $f:\mathcal{X}\rightarrow\Delta^{K}$ map inputs $\boldsymbol{x}_{i}\in\mathcal{X}$ to predicted probabilities $\hat{\boldsymbol{p}}_{i}=f(\boldsymbol{x}_{i})=[\hat{p}_{i1},\hat{p}_{i2},\ldots,\hat{p}_{iK}]^{\top}\in\Delta^{K}$. Let $\boldsymbol{y}_{i}=[y_{i1},y_{i2},\ldots,y_{iK}]^{\top}\in\Delta^{K}$ be the one-hot encoded true label. We define the Focal Calibration Loss as
@@
$\mathcal{L}_{\text{FCL}}^{\gamma,\lambda}\;=\;\frac{1}{N}\sum_{i=1}^{N}\!\Bigl{[}\,\mathcal{L}_{\text{focal}}\bigl{(}f(\boldsymbol{x}_{i}),\boldsymbol{y}_{i}\bigr{)}\;+\;\lambda\,\mathcal{L}_{\text{calib}}\bigl{(}f(\boldsymbol{x}_{i}),\boldsymbol{y}_{i}\bigr{)}\Bigr{]},$ (10)
@@
where $\gamma\geq 0$ and $\lambda\geq 0$ control the focal “focusing” and calibration trade-off, respectively. Specifically:
$\mathcal{L}_{\text{focal}}(f(\boldsymbol{x}_{i}),\boldsymbol{y}_{i})=-\sum_{k=1}^{K}\;y_{ik}\,\bigl{(}1-\hat{p}_{ik}\bigr{)}^{\gamma}\,\log\hat{p}_{ik},$ (11)
@@
$\mathcal{L}_{\text{calib}}(f(\boldsymbol{x}_{i}),\boldsymbol{y}_{i})=\left\|f(\boldsymbol{x}_{i})-\boldsymbol{y}_{i}\right\|_{2}^{2}=\sum_{k=1}^{K}\bigl{(}\hat{p}_{ik}\;-\;y_{ik}\bigr{)}^{2}.$ (12)
In practice, $\boldsymbol{y}_{i}$ is the (one-hot) observed class label for $\boldsymbol{x}_{i}$. By contrast, $\boldsymbol{\eta}(\boldsymbol{x})$ in certain theoretical derivations is the true conditional probability distribution $P(y\mid\boldsymbol{x})$, which is unknown in real-world datasets. Eq. (10) thus provides a fully empirical training objective and we define $\sigma^{\gamma,\lambda}(q):=(1-q)^{\gamma}-\gamma\,q\log(q)\,(1-q)^{\gamma-1}-2\lambda\,q,$ whose strict monotonicity and unique root properties are established in Appendix A.6.
Why FCL Helps.
Adding the squared-error calibration term $\lambda\,\|\hat{\boldsymbol{p}}_{i}-\boldsymbol{y}_{i}\|_{2}^{2}$ directly rewards probability estimates that approach the one-hot label, thereby nudging predictions toward lower calibration error. Even though one-hot labels are a crude proxy for $\boldsymbol{\eta}(\boldsymbol{x})$, this step ensures that, over many samples, the model’s output distribution more closely aligns with empirical label frequencies.
4 Theoretical Evidence
4.1 Mitigating OC/UC Beyond Focal Loss
Theorem 4.1 (Reducing Over-/Underconfidence).
Let $f$ be a classifier trained by minimizing $\mathcal{L}_{\text{FCL}}^{\gamma,\lambda}$ with $\lambda>0$. Suppose $\boldsymbol{\eta}(\boldsymbol{x})$ is the (unknown) true conditional probability vector. Then for any $\boldsymbol{x}\in\mathcal{X}$,
$\bigl{|}\,\max_{k}\hat{p}_{k}(\boldsymbol{x})\;-\;\max_{k}\eta_{k}(\boldsymbol{x})\bigr{|}\;\leq\;\|\,\hat{\boldsymbol{p}}(\boldsymbol{x})\;-\;\boldsymbol{\eta}(\boldsymbol{x})\|_{2}.$ (13)
Although $\boldsymbol{\eta}(\boldsymbol{x})$ is not directly accessible, this theorem shows that, if the model predictions $\hat{\boldsymbol{p}}(\boldsymbol{x})$ approach $\boldsymbol{\eta}(\boldsymbol{x})$ in $\ell_{2}$-distance, the gap between their respective maximum components shrinks. As a result, reducing $\|\hat{\boldsymbol{p}}(\boldsymbol{x})-\boldsymbol{\eta}(\boldsymbol{x})\|_{2}$ (which is encouraged by $\mathcal{L}_{\text{calib}}$) inherently bounds over-/underconfidence. Formal details in Appendix A.2 and  A.3.
4.2 $\mathcal{L}_{\text{FCL}}^{\gamma,\lambda}$ is Classification-Calibrated
Next, we show that minimizing the focal calibration risk $\mathcal{R}^{\mathcal{L}_{\text{FCL}}^{\gamma,\lambda}}$ achieves the Bayes-optimal classification boundary (Zhang, 2004). Although calibration term in Eq. (12) does not necessarily recover the exact posterior $\boldsymbol{\eta}(\boldsymbol{x})$, we prove it preserves the ranking of classes required for classification consistency (Bartlett et al., 2006; Tewari & Bartlett, 2007).
Definition 4.2 (Pointwise Conditional Risk).
Let $\mathcal{L}\colon\Delta^{K}\times\Delta^{K}\to\mathbb{R}$ be a surrogate loss, and let $\boldsymbol{e}_{y}\in\{0,1\}^{K}$ be the one-hot label for class $y$. For an input $\boldsymbol{x}\in\mathcal{X}$ with true class-posterior $\boldsymbol{\eta}(\boldsymbol{x})=[\eta_{1}(\boldsymbol{x}),\dots,\eta_{K}(\boldsymbol{x})]^{\top}$, and a predicted distribution $\boldsymbol{q}(\boldsymbol{x})\in\Delta^{K}$, the pointwise conditional risk is
$\mathcal{R}^{\mathcal{L}}\!\bigl{(}\boldsymbol{q}(\boldsymbol{x});\boldsymbol{\eta}(\boldsymbol{x})\bigr{)}\;=\;\sum_{y\in\mathcal{Y}}\eta_{y}(\boldsymbol{x})\,\mathcal{L}\bigl{(}\boldsymbol{q}(\boldsymbol{x}),\,\boldsymbol{e}_{y}\bigr{)}.$ (14)
Definition 4.3 (Classification-Calibrated (Bartlett et al., 2006; Tewari & Bartlett, 2007)).
A loss $\mathcal{L}$ is said to be classification-calibrated if its minimizer preserves the same $\arg\max$ ordering as the true posterior. Formally, let
$\boldsymbol{q}^{*}(\boldsymbol{x})=\arg\min_{\boldsymbol{q}\in\Delta^{K}}\mathcal{R}^{\mathcal{L}}\bigl{(}\boldsymbol{q}(\boldsymbol{x});\boldsymbol{\eta}(\boldsymbol{x})\bigr{)}.$ (15)
Then $\mathcal{L}$ is classification-calibrated if, for every $\boldsymbol{x}$, $\arg\max_{y}\,\bigl{[}q_{y}^{*}(\boldsymbol{x})\bigr{]}\;=\;\arg\max_{y}\,\bigl{[}\eta_{y}(\boldsymbol{x})\bigr{]}.$
Intuitively, if a surrogate loss is classification-calibrated, minimizing its pointwise risk yields a decision rule (via $\arg\max$) that matches the Bayes rule, hence achieving Bayes-optimal classification (Zhang, 2004). In Appendix A.5, we prove that $\mathcal{L}_{\text{FCL}}^{\gamma,\lambda}$ satisfies the strictly order-preserving property, ensuring $\arg\max_{y}\hat{p}_{y}(\boldsymbol{x})$ coincides with $\arg\max_{y}\eta_{y}(\boldsymbol{x})$. Consequently, $\mathcal{R}^{\mathcal{L}_{0\text{-}1}}\bigl{(}f^{\boldsymbol{q}^{\gamma,\lambda,*}}\bigr{)}\;=\;\mathcal{R}^{\mathcal{L}_{0\text{-}1}}\bigl{(}f^{\mathcal{L}_{0\text{-}1},*}\bigr{)},$ indicating that $\mathcal{L}_{\text{FCL}}^{\gamma,\lambda}$ recovers the Bayes-optimal decision boundary. Thus, by minimizing FCL, we obtain a classifier whose $\arg\max$ matches the optimal $\arg\max$ implied by $\boldsymbol{\eta}$, meeting the requirement for Def. 4.3.
4.3 Strictly Proper
Although classification calibration ensures correct $\arg\max$ decisions, many real-world tasks also require accurate probability estimates. A loss $\mathcal{L}$ is strictly proper if it is minimized only when $\hat{\boldsymbol{p}}(\boldsymbol{x})$ exactly matches the true probability $\boldsymbol{\eta}(\boldsymbol{x})$. While the true $\boldsymbol{\eta}(\boldsymbol{x})$ remains unknown, strictly proper losses incentivize unbiased probability estimation (Reid & Williamson, 2010; Williamson et al., 2016).
Definition 4.4 (Strictly Proper Loss).
$\mathcal{L}$ is strictly proper if, for any $\boldsymbol{y}\in\Delta^{K}$ and any $\hat{\boldsymbol{p}}\neq\boldsymbol{y}$, $\mathcal{L}(\boldsymbol{y},\boldsymbol{y})\;<\;\mathcal{L}(\hat{\boldsymbol{p}},\boldsymbol{y}).$
We show in Appendix A.7 that the calibration term $\|\hat{\boldsymbol{p}}-\boldsymbol{y}\|_{2}^{2}$ is strictly proper. We further prove in Appendix A.8 that FCL is strictly proper if $\lambda>0$. Intuitively, adding $\|\hat{\boldsymbol{p}}-\boldsymbol{y}\|_{2}^{2}$ penalizes deviations from the observed label distribution more severely as $\hat{\boldsymbol{p}}$ drifts away, forcing the model to converge to correct probability estimates in expectation.
4.4 Reduction of Post-Processing Gap
We finally show that FCL reduces the post-processing gap (Blasiok et al., 2024) relative to focal loss alone. Recall that post-processing gap measures how much a smooth recalibration function could further improve the model’s calibration:
Theorem 4.5.
@@ 4.4 Reduction of Post-Processing Gap
For any predictor $f:\mathcal{X}\to[0,1]$ and any distribution $\mathcal{D}$ over $\mathcal{X}\times\{0,1\}$, the $\mathcal{L}_{\text{FCL}}^{\gamma,\lambda}$ yields a smaller (or equal) post-processing gap than focal loss alone:
$\text{pGap}_{\mathcal{D}}\!\bigl{(}\mathcal{L}_{\text{FCL}}\bigr{)}\;\leq\;\text{pGap}_{\mathcal{D}}\!\bigl{(}\mathcal{L}_{\text{focal}}\bigr{)}.$ (16)
This implies that by incorporating $\|\hat{\boldsymbol{p}}-\boldsymbol{y}\|_{2}^{2}$ into training, the model’s raw outputs require less recalibration after the fact. Additional details and proofs are in Appendix A.14.
5 Experimental Results